OpenStudy (anonymous):

Rosie works 40 hours a week as an office manager. If she made $23,920 last year, how much was she paid per hour?

OpenStudy (anonymous):

NOT MY WORK THIS IS @TheStudyOwl Hi, Ok first we need to set up the problem. She made $23,920 last year. Well how many weeks are in a year? 365÷7=52 Ok so theres 52 weeks in a year. Now how much does she make in a week? 23,920÷52=460 Ok so she makes $460 in a week. Ok now how much money does she make in a day? 460÷7=65 Ok so she makes $65 in a day. Do you think you can finish it?(: Let me know if you still need help(: From this question http://openstudy.com/study#/updates/51abc007e4b08f9e59e378d9

OpenStudy (anonymous):

ok so if she made 65$ a day how much did she make an hour??

OpenStudy (anonymous):

24 hours in a day, take 65/24